# Cascade Speculative Drafting for Even Faster LLM Inference

### 16 Feb 2024 v3

@@ 3.3 Max-Gram for Better Statistical Drafting
As both Vertical Cascade and Horizontal Cascade remark cascade toward faster draft models, a statistical language model, which is the basis of the cascade, becomes essential for the efficiency of both approaches. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. In light of this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality). Surprisingly, despite its simplicity, this straightforward and general algorithm has proven highly effective in aligning with neural language model outputs. We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
@@ 3.4 Algorithm
The algorithm requires an upper-triangular hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the lenience parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
@@ 6.2 Speculative Decoding
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is proposed to improve inference speed by using a draft model to perform auto-regressive decoding and a target model to check and replace the generated tokens. A key to further speed up is to increase the acceptance rate in speculative decoding, which can be done by aligning the distribution between the draft model and the target model. Some recent work has proposed to improve the performance of speculative decoding in a domain/task-specific setting. Zhou et al. (2023) use generalized knowledge distillation to finetune the draft model and achieve 10%-45% speedup compared to draft models without knowledge distillation. Liu et al. (2023) propose Online Speculative Decoding to continually update multiple draft models on observed user query data, and the draft models align with the query distribution in real-time. However, both approaches require additional task-specific training making it difficult to apply the methods in a general setting. On the other hand, He et al. (2023) propose constructing a datastore from a domain-specific corpus to serve as a draft model which can be used when a draft model is not available.

### 27 Feb 2024 v4

@@ 3.3 Max-Gram for Better Statistical Drafting
As both Vertical Cascade and Horizontal Cascade remark cascade toward faster draft models, a statistical language model, which is the basis of the cascade, becomes essential for the efficiency of both approaches. In our pursuit of a more effective statistical language model, we noticed a general pattern: in language model generation, some words and phrases from the input query frequently reappear in the generated content. In light of this observation, we designed the Max-Gram (MaG) algorithm. It greedily identifies maximal matches between the initial input (or existing generation) and tokens from the end of the generation. In cases where there is no match, we resort to a bigram model based on the probability distribution of Wikipedia (chosen to maintain the generality). We include a GPU-friendly version of the Max-Gram algorithm in Appendix A.
@@ 3.4 Algorithm
The algorithm requires an upper-triangular hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the lenience $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of lenience without altering the output distribution.
@@ 6.2 Speculative Decoding
With the success of Speculative Decoding (Chen et al., 2023; Leviathan et al., 2023) in reducing the large language model inference latency, some recent works have attempted to improve Speculative Decoding by reducing the rejection rate. Zhou et al. (2023) propose using generalized knowledge distillation and achieve a lower rejection rate compared to other knowledge distillation methods. Avoiding an additional draft model, self-drafting is an approach to speculative decoding by reusing part of the target model together with added weight to perform drafting (Zhang et al., 2023; Hooper et al., 2024). Tree attention involves generating multiple candidates during drafting to increase the chance of acceptance (Spector & Re, 2023; Miao et al., 2024). Besides reducing the rejection rate, improving drafting efficiency can also reduce latency. Spector & Re (2023) propose using speculative decoding for drafting, showing similarities to the vertical cascade; however, their method only has two layers of speculative decoding and does not observe the recursive nature of the vertical cascade nor the lenience among draft models, two crucial aspects for the performance of vertical cascade.